OpenStudy (anonymous):

Some time ago you invested $10,000 at interest. You know the formula for money, M(t), in terms of time t is M(t)=10000(1.0325)^t where t is measured in years. When does the investment double? This is time from time zero, the day you made the investment.

Directrix (directrix):

At time zero, the amount is $10,000. The task is to find when the amount will be $20,000. M(t)=10000(1.0325)^t 20,000 = 10,000* (1.0325)^t 2 = (1.0325)^t Solve that equation for t. -->@theworldfines1 Post what you can and we can compare answers, okay?

OpenStudy (anonymous):

Thanks for the tip. This is what i came up with. (1.0325)^t=2 t In(1.0325)=In(2) t = In(2)/In(1.0325)

Directrix (directrix):

Crank out those natural logs to get a numerical value for t.

Directrix (directrix):

Post what you get.

Directrix (directrix):

@theworldfines1

OpenStudy (anonymous):

21.67